How many different ways can the letters in bird be arranged?
4!
its called four factorial
4 * 3 * 2 *1 = 24 ways
What are the chances of the moon crashing into earth?
The chances are actually not known exactly, but it is estimated a 10/90 chance. The two options are it either free's itself from earth's gravitational pull about a couple billion years from now or it is hit by some massive asteroid and screws up its orbit causing it to come crashing or getting freed from, once again, earths gravitational pull. It is more likely that it will just float away a couple billion years from now.
The ratio tells us that for every 5 red marbles there are 4 blue ones.
Now the total number is 225.
Out of 9 total marbles, we know 4 are blue.
So we can look at 4 blue/9 total= (x blues which we want to know)/ 225.
This is not hard to solve by cross multiplication.
9x=4(225)
or 9x=900 so x is 100 and there are 100 blue marbles.
Now if that is so, we know the rest of the marbles must be red so there are 225-100=125 red ones.
Let us check that this ratio is 5/4 as desired.
Red/Blue=125/100 and 25 goes into 125 5 times and into 100 4 times so this becomes
5/4.
How else would we do this problem. Since the ratio of 5/4, we know as we said above there are 4/9 blue marbles ( every 4 marbles out of 9 is blue). Now 4/9(225)=4x225/9=100.
What is the probability of getting a sum of 7 when tossing 2 six sided dice?
To determine the probability of getting any number with 2 six sided dice you need two pieces of information
First consider all the possible ways to achieve a sum of 7.
7 is the most common roll based on probability. You can achieve it 6 different ways.
There are 36 possible number combinations when you roll 2 normal 6 sided die.
6/36 is the probability of rolling a 7. Simplify this fraction to 1/6.
So out of every 6 rolls there should be at least one 7 rolled based on normal probability.
81
What is the probability of drawing a queen and heart from a standard deck of cards?
The probability of drawing the queen of hearts is 1 in 52, or about 0.01923.
What is the probability of tossing a coin three times and getting it on heads each time?
Each toss has a 1/2 probability of getting heads. Each toss is an independent event. So three heads in a row (heads AND heads AND heads) would have a probability of:
1/2 * 1/2 * 1/2 = (1/2)^3 = 1/(2^3) = 1/8 = 12.5%
There are two types of roulette wheels, one with a 0 and another with a 0 AND 00.In the first case, the probability will be 0.324 and in the second it will be 0.316, approx.
What is 200000 increased to 6 percent?
OMG more easy... (Without equations)
percent increase = [amount * (1 + x%)]
x% = x/100; 6% its (6/100) = .06
200000*1.06 = 212,000
How many cards with a n odd number in a standard deck of cards?
Not counting the face cards nor Aces which do not have a number, there are 16 odd numbered cards.
When one card is drawn from a standard deck what is the probability of drawing a king?
Well, if there are 52 cards in the deck, the probability will be 4/52, which is 1/13. If there are Jokers in the deck, there will be 54 cards, so the probability will be 4/54, which is 2/27.
Hope this helps!
Can you still have twins without having them run in your family?
doubtful
Yes, only fraternal twinning (caused by releasing more than one egg) has a genetic connection, and that connection only holds on the mother's side. So, many twins are born with no family history of twins at all.
What does experimental probability mean in math?
The probability of an event may be measured experimentally or theoretically. In experimental probability, an experiment is conducted repeatedly. The probability of the event is the number of experiments in which the event occurs as a proportion of the number of times the experiment is conducted. By contrast, the theoretical probability is calculated from theoretical models and laws of science (and some assumptions about unbiased/fairness).
What is the probability of being dealt 4 consecutive face cards from a standard deck of cards?
The probability that the first four cards are face cards is (16/52)*(15/51)*(14/50)*(13/49) = 43680/6497400 = approx 0.0067
When two dice are rolled what is the probability of not getting a sum of 12?
The probability of rolling 12 with 2 dice is 1 in 36.
The probability of not rolling 12 with 2 dice is 35 in 36.
How many pints are there in 216 quarts?
Since there are 2 pints in 1 quart, there are 432 pints in 216 quarts.
Use the context: if the variable should be greater than the mean then z is positive and if less than the mean, it should be negative.
Is an event with a probability of 1 more or less likely than an event with a probability of 2?
It is more likely because it can exist. An event with a probability of 2 cannot exist.